An aqueous solution containing 17.3 g of an unknown molecular (nonelectrolyte) compound in 102.0 g of water was found to have a freezing point of -1.6 ?C.
Part A
Calculate the molar mass of the unknown compound.
Answer #1

Quantity of solute dissolved in 1kg of solvent = 17.3*1000/102= 169.6g

Freezing point depression constant for water =

1.86 degrees/mol

Given depression point = 1.6 degrees

Molar mass of solute = 169.6*1.86/1.6 = 197.16g/mol

Molar mass of unknown compound = 197.16g/mol